OpenStudy (anonymous):

The formula for the surface area of a cylinder is A=2pr(r+h).Find h if A=80p and r=5. 1 3 5 6

OpenStudy (anonymous):

80p=10p(5+h) 8=5+h h=3

OpenStudy (anonymous):

thank you i was making sure my answer was correct.

OpenStudy (anonymous):

yw